OpenStudy (anonymous):

write 3x + 4y = 8 in slope intercept form.

OpenStudy (anonymous):

i believe it is y = (-3/4)x + 2

OpenStudy (anonymous):

3x + 4y = 8 4y = -3x + 8 y = -3/4x + 2 you are correct :)

OpenStudy (anonymous):

thanks!